# A Component-Based Survey of Interactions between Large Language Models and Multi-Armed Bandits

### v1

@@ 1 Introduction
The organization of this article is as follows. Section 2 introduces the background information about large language models and multi-armed bandits. A component-based taxonomic framework for both LLM systems and bandit systems is presented in Section 4. We then review representative studies on the bidirectional integration between LLMs and bandit algorithms in Sections 5 and 6. Finally, we discuss challenges and future opportunities in Section 7 and conclude in Section 8.
@@ 5 Bandit-Based Enhancements for LLM Systems
This section provides a component-level review of how bandit algorithms are integrated into large language models to improve training, inference, and adaptation. Table LABEL:tab:bandit-for-llm offers a consolidated overview of selected works discussed in this section.
@@
Pre-training Static masking and curriculum heuristics fail under evolving data distributions. Adaptive selection of masking or training strategies using loss-based feedback. [79]
Multi-task pre-training requires balancing correlated tasks with heterogeneous learning dynamics. Structured task-level exploration and scheduling under uncertainty. [126]
Fixed data mixtures struggle to balance data quality, diversity, and domain shift at scale. Online adjustment of domain-level sampling ratios to optimize information gain. [3, 150, 72]
@@
Fine-tuning Static data sampling and reward signals cause reward overfitting and unstable updates. Adaptive reweighting and selection of training data under noisy feedback. [163, 162, 159, 155]
Preference annotation is costly, making random or static sampling inefficient. Active selection of informative preference queries and context segments. [33, 68, 118]
Limited compute budgets require allocating fine-tuning effort in non-stationary settings. Online allocation of fine-tuning budgets across models or datasets. [111, 142, 141]
Alignment Human preference feedback is sparse, noisy, and expensive to collect. Adaptive selection of informative preference queries under feedback uncertainty. [69, 73, 61]
Alignment involves iterative preference optimization with non-stationary feedback. Bandit-style exploration to balance exploitation and exploration during online alignment. [11, 18, 81]
Alignment pipelines include structured decisions beyond direct policy updates. Modeling structured input or context selection as bandit problems. [33]
Prompt Design & Selection Prompt performance is highly sensitive to prompt choices, while evaluation is costly and black-box. Modeling prompt variants as arms to enable budget-aware exploration and best prompt identification. [110, 109, 6]
Numeric rewards are unavailable or unreliable; feedback often comes as pairwise preferences or implicit comparisons. Preference-based and dueling bandits for prompt optimization under comparison-only feedback. [60, 136]
In-context example selection and ordering has combinatorial structure that heuristics handle poorly under limited trials. Bandit-based selection and ordering of in-context examples or trajectory prompts to improve in-context learning. [137, 106, 52]
@@
Large prompt spaces and privacy constraints require scalable optimization from logged or distributed feedback. Offline, neural, and federated bandit formulations for scalable prompt optimization under query and privacy limits. [51, 50, 119, 62, 70]
@@
Tool and Function Calling Multi-step tool use yields delayed and noisy execution feedback. Bandit-guided credit assignment for robust multi-step tool-use reasoning. [147, 22]
Large and evolving tool sets challenge generalization of static selection policies. Semantic abstraction and contextual representations for scalable tool selection. [80, 20]
Tool use in agentic settings involves long-horizon interactions and delayed rewards. Bandit and RL-based decision-making for long-horizon tool orchestration. [46, 148]
@@
Context Understanding LLMs must explore and adapt under rich context with sparse or costly feedback. Bandit-driven in-context exploration and feedback-efficient adaptation. [53, 104, 34, 120]
Long or structured inputs exceed context limits, making it unclear which segments matter. Adaptive context attribution/selection by treating chunks or passages as bandit arms. [33, 89]
Inference-time decisions must adapt to non-stationary user/task contexts, including safety-critical querying. Contextual bandits for online strategy/model selection and controlled query exploration. [99, 85, 97, 105]
Retrieval-Augmented Generation (RAG) Retrieval control must adapt to varying query complexity, context needs, and resource budgets. Bandit-based selection of retrieval strategies and intensity (e.g., retrieval ratio, prompt length, database partition). [121, 87, 132]
Relevant evidence is heterogeneous and context-dependent, making source/document selection unreliable under limited trials. Bandit-driven selection of knowledge sources, document subsets, and prompt formats conditioned on the query context. [152, 89]
RAG optimization is multi-objective and often involves hierarchical pipeline/configuration choices in non-stationary settings. Multi-objective and hierarchical bandit control to jointly optimize retrieval–generation trade-offs and tune pipeline variants/hyperparameters online. [122, 36]
Inference Optimization Inference-time routing/model selection must trade off output quality against monetary cost and system load, often under explicit budget or latency constraints that vary across queries and time. Bandit-based (contextual/preference-conditioned) routing to adaptively select models/backends using online performance–cost feedback. [57, 90, 124, 133]
Redundant computation across repeated or similar queries increases inference cost in large-scale deployment. Combinatorial bandit formulations for cache selection and reuse to reduce cost while preserving performance guarantees. [144]
@@
Decoding & Reranking Static decoding or reranking heuristics fail to generalize. Adaptive selection of decoding configurations using online feedback. [43, 65]
Speculative decoding provides richer-than-bandit verification signals. Exploiting partial or near full-information feedback for faster adaptation. [64]
Token-level decoding induces irreversible decision dependencies. Bandit modeling of token-level decisions to characterize achievable regret. [112]
@@
Adaptation & Personalization User preferences are heterogeneous and time-varying. Online exploration–exploitation of user-specific behaviors. [24, 38, 77, 76]
@@
Multi-task personalization must balance performance and cost. Bandit-based model selection and routing under user-specific trade-offs. [27, 57, 133]
@@
Inference-time personalization must adapt without retraining. Bandit-guided adaptation of reasoning or decoding trajectories. [112, 154]
@@ 5.3 Alignment
5.4 Prompt Design & Selection
@@ 6 LLM-Based Enhancements for Bandit Systems
We will systematically examine how LLMs enhance various components of multi-armed bandits. Some representative papers for each components are shown in Table LABEL:tab:llm-for-bandit.
@@
Regret Minimization Objective Heuristic or uniform exploration causes redundant sampling and slow convergence in large action spaces. Providing semantic or information-theoretic guidance to bias exploration toward informative regions. [113]
Classical regret ignores rich contextual priors available at decision time. Injecting contextual knowledge to reshape effective regret and reduce unnecessary exploration. [138]
Standard regret formulations overlook adaptive agent behavior over long horizons. Summarizing interaction history to guide regret-aware policy updates. [91]
Fixed regret objectives degrade under non-stationary rewards. Dynamically adjusting regret objectives using contextual and semantic signals. [28]
Arm Definition Discrete or unstructured arms fail to capture semantic relations, causing redundant exploration. Constructing semantically coherent arm representations via clustering or embedding. [128, 32]
Fixed arm definitions become suboptimal under evolving contexts or preferences. Dynamically redefining or personalizing arms through contextual reasoning. [138, 115]
Traditional arms cannot express structured or attribute-driven actions. Generating structured arms by extracting attributes and defining semantic clusters. [49]
Environment Stationary or low-dimensional assumptions fail in non-stationary or semantic-rich settings. Interpreting evolving contexts to guide adaptation and detect regime shifts. [28, 129, 2, 107]
Realistic environments are difficult to simulate for bandit training. Generating structured synthetic environments with diverse semantic dynamics. [49, 146]
Unstructured logs cannot be directly used as bandit context. Transforming free-form trajectories into structured contextual inputs. [5, 84]
Reward Formulation Static reward designs fail under non-stationary or ambiguous feedback. Adapting reward structures online via contextual interpretation. [35, 91]
Parametric rewards struggle with sparse or nonlinear feedback. Acting as reward surrogates through in-context regression or priors. [116, 84]
Human objectives are hard to encode as numeric rewards. Translating language specifications into explicit or multi-objective rewards. [13, 129]
Preference feedback is qualitative or pairwise. Converting language-based preferences into structured reward signals. [21, 139]
Sampling Strategy Naive exploration is costly in high-dimensional textual action spaces. Guiding information- and cost-aware exploration using semantic relevance. [113, 40, 19, 2, 27, 116]
Classical sampling rules generalize poorly across heterogeneous tasks. Learning meta-level exploration policies from textual interaction histories. [21, 108]
Explicit posteriors are hard to define for language-based rewards. Using natural language as an implicit posterior to guide exploration. [41, 58, 5]
hhahahah Action Decision Hand-crafted rules struggle with high-dimensional context and semantic feedback. Refining value or preference estimates via semantic encoding before decision rules. [10, 130, 40, 67, 151, 116]
Explicit decision formulas are brittle with sparse statistics. Using LLMs directly as policies mapping language summaries to actions. [21, 41, 58]
Language-based decisions may violate confidence principles. Constraining LLM decisions with posterior sampling or confidence bounds. [108, 5]

### v2

@@ 1 Introduction
The organization of this article is as follows. Section 2 introduces the background information about large language models and multi-armed bandits. Section 3 describes the survey methodology adopted in this work. A component-based taxonomic framework for both LLM systems and bandit systems is presented in Section 4. We then review representative studies on the bidirectional integration between LLMs and bandit algorithms in Sections 5 and 6. Finally, we discuss challenges and future opportunities in Section 7 and conclude in Section 8.
@@ 5 Bandit-Based Enhancements for LLM Systems
This section provides a component-level review of how bandit algorithms are integrated into large language models to improve training, inference, and adaptation. Table 3 offers a consolidated overview of selected works discussed in this section.
@@
Pre-training Static masking strategy selection Bandit-based dynamic masking pattern selection [79]
Static task sampling in multi-task pre-training Structured multi-task bandit formulation [126]
Suboptimal trade-off between data quality and diversity Online domain-level sampling ratio optimization [3, 150, 72]
@@
Fine-tuning Reward overfitting in RL-based fine-tuning Adaptive training data reweighting and selection [163, 162, 159, 155]
Inefficient preference data collection Bandit-based preference and context selection [33, 68, 118]
Static budget allocation under non-stationarity Bandit-based adaptive dataset mixture optimization [111, 142, 141]
Alignment Costly and inefficient preference feedback collection Adaptive bandit-based preference query selection [69, 73, 61]
Non-stationary and myopic alignment dynamics Bandit-style exploration in online preference learning [11, 18, 81]
Structured decision stages in alignment pipelines Bandit-guided structured alignment pipelines [33]
Prompt Design and Selection Budget-limited prompt evaluation Bandit-based best prompt identification [110, 109, 6]
Implicit and costly prompt feedback Preference-based and dueling bandit optimization. [60, 136]
Suboptimal prompt composition Bandit-based example and trajectory selection [137, 106, 52]
@@
Large-scale and privacy-constrained prompt spaces Offline, neural, and federated bandit optimization [51, 50, 119, 62, 70]
@@
Tool and Function Calling Delayed and noisy feedback in multi-step tool use Execution-feedback-driven bandit tool selection [147, 22]
Large and heterogeneous tool spaces Semantic-aware bandit-based tool generalization [80, 20]
Delayed feedback and credit assignment Bandit-informed agentic tool optimization [46, 148]
@@
Context Understanding Exploration and adaptation under sparse feedback Bandit-based in-context exploration [53, 104, 34, 120]
Inefficient long-context utilization Adaptive bandit-guided context selection [33, 89]
Non-stationary inference-time decisions Contextual bandit-based online strategy selection. [99, 85, 97, 105]
Retrieval-Augmented Generation (RAG) Adaptivity requirements in retrieval control across queries and budgets Bandit-based retrieval strategy selection [121, 87, 132]
Heterogeneity and context dependence of retrieved evidence Bandit-driven knowledge source selection [152, 89]
Multi-objective and hierarchical RAG optimization. Hierarchical and multi-objective bandit optimization [122, 36]
Inference Optimization Static and inefficient inference model selection Adaptive bandit-based LLM routing [57, 90, 124, 133]
High inference cost and resource waste Bandit-driven cache and resource optimization [144]
@@
Decoding and Reranking Limited generalization of static decoding and reranking heuristics Adaptive decoding configuration selection [43, 65]
Rich verification signals in speculative decoding Verification-augmented online decoding optimization [64]
Irreversible dependencies in token-level decoding Token-level bandit modeling [112]
@@
Adaptation and Personalization Heterogeneity and non-stationarity of user preferences Online user behavior exploration–exploitation [24, 38, 77, 76]
@@
Performance–cost trade-offs in multi-task personalization Bandit-based personalized model selection and routing [27, 57, 133]
@@
Inference-time personalization without retraining Bandit-driven inference-time preference adaptation [112, 154]
@@ 5.1 Pre-training
Despite these advances, existing bandit-based pre-training methods still face limitations in reward design and granularity. Most approaches rely on short-horizon or proxy-based rewards (e.g., loss or perplexity) and operate at coarse domain or task levels, leaving open challenges in aligning bandit feedback with long-term generalization and in extending adaptive control to finer-grained data or training decisions.
@@ 5.3 Alignment
5.4 Prompt Design and Selection
@@ 6 LLM-Based Enhancements for Bandit Systems
This section systematically examines how large language models (LLMs) enhance various components of multi-armed bandit algorithms. Table 4 presents a consolidated overview of representative papers corresponding to each component.
@@
Regret Minimization Objective Slow convergence under uniform regret minimization LLM-guided information–regret trade-off optimization [113]
Costly exploration in real-time decision scenarios Contextual-prior–driven regret minimization [138]
Static regret objectives under learning dynamics Regret-centric adaptive agent optimization [91]
Regret degradation under shifting rewards Dynamic regret objective adjustment [28]
Arm Definition Redundant exploration in high-dimensional action spaces LLM-based semantic arm compression [128, 32]
Static arms under preference drift Dynamically redefining or personalizing arms through contextual reasoning [138, 115]
Limited expressiveness of discrete arms LLM-generated structured arm manifolds [49]
Environment Stationary or low-dimensional assumptions fail in non-stationary or semantic-rich settings Interpreting evolving contexts to guide adaptation and detect regime shifts. [28, 129, 2, 107]
Realistic environments are difficult to simulate for bandit training Generating structured synthetic environments with diverse semantic dynamics [49, 146]
Unstructured and latent environment states Language-to-context environment mapping [5, 84]
Reward Formulation Noisy and ambiguous feedback Context-aware dynamic reward shaping [35, 91]
Sparse or ethically constrained rewards LLM-based reward prior and regression [116, 84]
Misalignment between numeric rewards and human goals Natural-language-to-reward translation [13, 129]
Long-horizon credit assignment difficulty LLM-mediated preference and algorithmic rewards [21, 139]
Sampling Strategy Naive exploration is costly in high-dimensional textual action spaces LLM-guided informative sampling [113, 40, 19, 2, 27, 116]
Classical sampling rules generalize poorly across heterogeneous tasks Learning meta-level exploration policies from textual interaction histories [21, 108]
Explicit posteriors are hard to define for language-based rewards Using natural language as an implicit posterior to guide exploration [41, 58, 5]
Action Decision Poor action ranking under complex context LLM-enhanced value and preference estimation [10, 130, 40, 67, 151, 116]
Rigid numerical decision heuristics LLM-as-policy action selection [21, 41, 58]
Inconsistent confidence-based decisions LLM-aligned posterior-driven action choice [108, 5]